# Market-Driven Equilibria for Distributed Solar Panel Investment

### v1

@@ VII-B1 Validity of Theorem 2
Given this setup, we first compute the expected revenues of panel investors under the real-time markets using (25) and (26). Meanwhile, for the contract-based market, we obtain the aggregate demand function, i.e., $\widehat{d}^{\star}(\pi)$, using (27). Leveraging the characterizing equations for the heterogeneous case, we then obtain the NE aggregate investment capacity under the three considered short-term markets as summarized in Table II.
@@
First, observe that this result mirrors the comparison outcome established in (20) for case 1 (general case) in Theorem 1, and likewise in Theorem 2. Moreover, although in this setup we do not scale the distribution of consumers’ premiums, i.e., $\epsilon=1$ is not in the neighborhood of zero, and assumption (30) does not hold, we still obtain the same comparison result as stated in (22), corresponding to case 2 (small solar premium) of Theorem 1, and likewise in Theorem 2. Specifically, the relation $c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}}$ in that case holds here with a strict inequality, i.e., $c_{\mathrm{opt}}<c^{\mathrm{ne}}_{\mathrm{cb}}$. We then consider the scenario where consumers assign no premium to solar energy, i.e., $\epsilon=0$, which yields identical aggregate capacities across all market mechanisms, as shown in Table II and implied by (24) under case 3 (no solar premium) of Theorems 1 and 2.
In the following two subsections, we analyze how variations in consumers’ solar premiums and in panel capital and installation costs affect the comparison of equilibrium aggregate capacities across market mechanisms.
@@ VII-B2 Sensitivity analysis with respect to consumers’ solar premiums
As consumers’ solar premiums increase, this figure reveals two key insights. First, the single-product real-time energy market results in a larger gap compared to the social welfare optimal capacity. This is primarily due to this market’s inability to capture the additional revenue potential from environmentally conscious consumers who are willing to pay a premium for clean energy. Second, the contract-based market consistently results in over-investment, reflected in the widening gap between $c^{\mathrm{ne}}_{\mathrm{prt}}$ and $c^{\mathrm{ne}}_{\mathrm{cb}}$.
Figure 4: Variation of equilibrium solar panel investments in response to consumers’ solar premium scaling
@@ VII-B3 Sensitivity analysis with respect to capital and installation costs
We vary the capital and installation costs for a unit panel capacity and analyze how these changes affect the equilibrium aggregate capacities, as shown in Fig. 5.
Figure 5: Equilibrium solar panel variation with changes in installation costs

### v2

@@ VII-B1 Validity of Theorem 2
Given this setup, we first compute panel investors’ expected revenues under the real-time markets using (27) and (28). For the contract-based market, we first obtain the aggregate demand function, $\widehat{d}^{\star}(\pi)$, from (30) and then compute investors’ revenue. Using the characterizing equations for the heterogeneous case, we derive the NE aggregate investment capacity under the three short-term markets, as summarized in Table II.
@@
First, observe that this result mirrors the comparison outcome established in (22) for case 1 (general case) in Theorem 1, and likewise in Theorem 2. Moreover, although in this setup we do not scale the distribution of consumers’ premiums, i.e., $\epsilon=1$ is not in the neighborhood of zero, and assumption (33) does not hold, we still obtain the same comparison result as stated in (24), corresponding to case 2 (small solar premium) of Theorem 1, and likewise in Theorem 2. Specifically, the relation $c_{\mathrm{opt}}\lesssim c^{\mathrm{ne}}_{\mathrm{cb}}$ in that case holds here with a strict inequality, i.e., $c_{\mathrm{opt}}<c^{\mathrm{ne}}_{\mathrm{cb}}$. We then consider the scenario where consumers assign no premium to solar energy, i.e., $\epsilon=0$, which yields identical aggregate capacities across all market mechanisms, as shown in Table II and implied by (26) under case 3 (no solar premium) of Theorems 1 and 2. As Table II suggests, discrepancies in equilibrium aggregate capacities are modest in the base case ($\epsilon=1$) because solar premiums are small on average (2.86 ¢\$/kWh) relative to the utility price (29 ¢\$/kWh). As premiums increase, these discrepancies are expected to widen, as analyzed next.
In the following two subsections, we analyze how variations in consumers’ solar premiums and in panel capital and installation costs affect the equilibrium aggregate capacities.
@@ VII-B2 Sensitivity analysis with respect to consumers’ solar premiums
As consumers’ solar premiums increase, this figure reveals two key insights. First, the single-product real-time market results in a larger gap compared to the social welfare optimal capacity. This is primarily due to this market’s inability to capture the additional revenue potential from environmentally conscious consumers. Second, the contract-based market consistently results in over-investment, reflected in the widening gap between $c^{\mathrm{ne}}_{\mathrm{prt}}$ and $c^{\mathrm{ne}}_{\mathrm{cb}}$.
Figure 4: Variations of equilibrium aggregate panel investment capacities in response to consumers’ solar premium scaling
@@ VII-B3 Sensitivity analysis with respect to capital and installation costs
We vary the capital and installation costs for a unit panel capacity and analyze how these changes affect the equilibrium aggregate panel investment capacities, as shown in Fig. 5.
Figure 5: Equilibrium aggregate panel investment capacities variation in response to changes in installation and capital costs